# Factorized Video Generation: Decoupling Scene Construction and Temporal Synthesis in Text-to-Video Diffusion Models

### v1

@@ Abstract
State-of-the-art Text-to-Video (T2V) diffusion models can generate visually impressive results, yet they still frequently fail to compose complex scenes or follow logical temporal instructions. In this paper, we argue that many errors, including apparent motion failures, originate from the model’s inability to construct a semantically correct or logically consistent initial frame. We introduce Factorized Video Generation (FVG), a pipeline that decouples these tasks by decomposing the Text-to-Video generation into three specialized stages: (1) Reasoning, where a Large Language Model (LLM) rewrites the video prompt to describe only the initial scene, resolving temporal ambiguities; (2) Composition, where a Text-to-Image (T2I) model synthesizes a high-quality, compositionally-correct anchor frame from this new prompt; and (3) Temporal Synthesis, where a video model, finetuned to understand this anchor, focuses its entire capacity on animating the scene and following the prompt. Our decomposed approach sets a new state-of-the-art on the T2V CompBench benchmark and significantly improves all tested models on VBench2. Furthermore, we show that visual anchoring allows us to cut the number of sampling steps by 70% without any loss in performance, leading to a substantial speed-up in sampling. Factorized Video Generation offers a simple yet practical path toward more efficient, robust, and controllable video synthesis. Project website: https://vita-epfl.github.io/FVG/.
@@ 1 Introduction
Figure 1: Example failure modes in SoTA video generative foundational models. We compare the first frame of videos from T2V Wan 2.2 [wan2025wan] and our Factorized Wan $2.$ T2V Wan 2.2 composes scenes incorrectly and exhibits logical temporal inconsistencies, struggling to establish coherent scene structure without explicit visual grounding. This can be solved via our factorization.
@@
Our analysis across multiple state-of-the-art systems reveals that many errors originate in the model’s representation of the scene itself. In Figure 1, we observe that existing models routinely (1) miscompose visual elements of the scene, and (2) generate initial states incompatible with the requested temporal evolution. Once the scene composition violates the prompt semantics, the temporal modeling fails. For example, if the model wrongly understands the context by assuming a “chicken” as a “roasted chicken” (see 1st sample), it becomes impossible to satisfy the prompt “a chicken tasting the sauce,” regardless of the strength of the motion model. This insight indicates a critical gap: Current T2V models may have over-entangled the two distinct tasks, static scene construction and temporal synthesis, yet are not structurally equipped to handle the first reliably.
@@
To delve into and understand this gap, we explore a diagnostic study by comparing Text-to-Video (T2V), Image-to-Video without text prompts (I2V), and Image-to-Video with text prompts (I2V+text) models in Table 1, revealing two observations. T2V models frequently produce distributionally shifted frames yet still achieve FVD scores comparable to I2V models, indicating that their motion modeling remains reasonably natural even when scene fidelity is relatively poor. I2V models exhibit the complementary behavior, strong FID from accurate initial scenes and weaker temporal coherence, while I2V+text balances both aspects. This contrast suggests a structural mismatch in current T2V models: scene grounding and temporal synthesis benefit from distinct inductive biases, yet existing architectures attempt to learn both simultaneously within a single model.
@@
These observations motivate us to revisit T2V. Rather than formulating it as a single task, we factorize it into two fundamentally factorized sub-problems: specifying the initial scene and generating its temporal evolution. A natural way to operationalize this separation is through a visual anchor defining the intended starting state of the video. However, such an anchor cannot be generated by simply running a Text-to-Image (T2I) model on the original video prompt, which often contains multi-stage actions or temporal transformations. Obtaining an appropriate anchor requires first clarifying the prompt’s intended initial condition.
@@
Based on this insight, we propose Factorized Video Generation (FVG), a pipeline that explicitly decomposes T2V generation into specialized stages designed to address each subproblem independently: (1) Reasoning: A Large Language Model (LLM) rewrites the video prompt into an explicit description of the initial scene, removing temporal transitions and eliminating logical ambiguities. (2) Composition: A T2I model uses this refined prompt to synthesize a high-quality, semantically grounded anchor frame. (3) Temporal synthesis: A video diffusion model conditions on both the anchor and the original prompt, focusing exclusively on generating coherent temporal dynamics. The contributions of this work are as follows:
@@
  • We introduce FVG, a simple and modular factorization of T2V generation into reasoning, composition, and animation, enabling improved scene grounding and temporal coherence.
@@
We also release our models, code, benchmark prompts and image anchors generations to support reproducibility and facilitate further research.
@@ 2 Related Work
Spatial-Temporal Factorization. To address the complexity in video-centric learning, factorization [tran2018closer] has become a fundamental architectural strategy [liu2022video, bertasius2021space, lin2019tsm, ma2024latte], which decouples spatial and temporal components to enable efficient video representation. Some works [chen2023videocrafter1, videoworldsimulators2024] extend this paradigm via hierarchical factorization across multiple scales. Due to the reduced model scale, these factorized approaches compromise spatial-temporal coupling, leading to suboptimal results when data is scaled [ma2024latte]. Differently, we propose a new perspective by lifting the architecture factorization to a higher task-oriented level, thereby preserving model capacity while achieving superior text-following capabilities in video generation.
Figure 2: Overview of the finetuning and inference pipelines for our factorized video generation. Anchor-Grounding Finetuning: The Video Diffusion Model (VDM) is trained to follow a visual anchor by injecting a clean image latent at a randomly chosen frame position and setting its diffusion timestep to $t=0$. Lightweight LoRA finetuning trains the model to treat this clean frame as a fixed scene constraint that guides the rest of the video. Factorized Inference Pipeline: a LLM modifies the video prompt into a first-frame descriptive prompt to generate an anchor image. The anchor is then injected into the VDM with the timestep fixed set to 0.
@@ 4 Methods
The goal of this work is to investigate how grounding text-to-video (T2V) diffusion models with at least one visual anchor affects their performance on complex compositional tasks. By introducing a visual anchor, we aim to decouple video generation into two subtasks, scene context construction and temporal motion modeling, and evaluate whether this separation improves the model’s understanding of scenes, objects, and their interactions. Our focus extends beyond visual quality to assess reasoning-oriented capabilities of video diffusion models. To this end, we compare standard T2V models with their text–image–to–video (I2V+text) counterparts across benchmarks designed to evaluate compositional video generation. Furthermore, to enable a fair comparison within the same architecture, we propose dedicated finetuning and inference pipelines that integrate a visual anchor into a pure T2V model. The following section details these pipelines which are summarized in Fig. 2.
@@ 4.1 Anchor-Grounding Finetuning
A core capability required by the FVG pipeline is the ability to inject a visual anchor into a pretrained T2V diffusion model. We enable this through anchor-grounding finetuning, which teaches the model to treat the provided image as the initial frame of the generated video.
@@ 4.1.2 Training Procedure
The training procedure for adapting a T2V model to an a visually anchored model is detailed in Algorithm 1. The core idea is to create a hybrid input $(z_{t},t_{vec})$ that pairs T1T-1 noisy frames at timestep $t$ with one clean anchor frame at $t=0$. The model $\epsilon_{\theta}$ is then finetuned via LoRA to denoise this out-of-distribution input, learning to treat the $t=0$ frame as a ground-truth anchor. The model itself is still conditioned only on the text prompt $y$.
@@
4.2 Factorized Inference Pipeline
At inference time, we generate a visual anchor from the input text prompt before initiating the video generation process. Given a video text prompt $y$, we use a large language model (LLM) to generate a plausible detailed description of the scene’s initial state $y_{0}$, which emphasizes spatial layout, appearance, and object configuration. This reformulated prompt $y_{0}$ is then passed to a pretrained image generation model $\mathcal{G}_{\text{img}}$ to synthesize the anchor image $I_{a}=\mathcal{G}_{\text{img}}(y_{0})$.
@@ 5.2 Performance on Complex Tasks
Results on T2V-CompBench: As shown in Tab. 2 across all models, adding an anchor image consistently improves compositional performances. All smaller Factorized models (CogVideo 5B, Wan 5B and Wan 1B) outperform the larger Wan 14B T2V model. Our factorized Wan 5B also outperforms commercial PixVerse-V3 baseline which is the best reported model on the benchmark. This demonstrates that visual grounding substantially enhances scene and action understanding even in smaller-capacity models. Within each model family, the factorized version outperforms the original model. Notably, our lightweight anchor-grounded LoRA on WAN 14B reaches performance comparable its pretrained I2V 14B variant (0.661 vs. 0.666), despite requiring no full retraining.
@@
Factorized I2V + Text 0.927 0.217 0.680 0.320 0.850 0.886 0.781 0.666 (+21.09%)
Factorized T2V (Ours) 0.929 0.183 0.696 0.328 0.839 0.891 0.757 0.661 (+20.18%)
@@
Factorized T2V (Ours) 0.920 0.384 0.656 0.262 0.790 0.844 0.402 0.608 (+32.75%)
Results on VBench2.0: We summarize in Tab. 3 the results on VBench 2.0, reporting the five aggregated metrics. The complete set of 18 underlying metrics is provided in the Appendix. We report only the composition score instead of the original creativity metric (the latter averages composition and diversity) and study the diversity specifically in Sec. 5.4. Across all architectures, factorizing T2V consistently improves every VBench metric except human fidelity, which shows a small decline even when using prompt upsampling. Similar to the observations on T2V-CompBench, our factorized Wan 5B outperforms the larger Wan 14B model, demonstrating that visual grounding is sometimes more beneficial than scaling. While performance gains are consistent, improvements on VBench are generally smaller than on T2V-CompBench. This is expected since VBench applies a stricter evaluation protocol as most metrics are binary (0 or 1 per video) with no partial credit. Nonetheless, the trend confirms that visual anchoring systematically enhances compositional reasoning and controllability even under stricter assessment criteria.
@@
Factorized T2V (Ours) 0.649 (+39.4%) 0.636 (+25.6%) 0.304 (+69.3%) 0.652 (–14.0%) 0.562 (+25.8%) 0.561 (+19.0%)
Qualitative Results.
Figure 5 shows representative examples comparing text-only generation with our factorized approach. Anchored videos consistently exhibit more accurate scene composition, stronger object–attribute binding, and clearer temporal progression.
Figure 3: Qualitative results showing our factorized method leads to better performance. Additional results are provided in the appendix.
@@ 5.3 Robustness to Reduced Sampling Steps
We assess robustness to reduced diffusion steps by decreasing sampling from 50 to 30 and 15 for Wan5B and calculate the results on T2V Compbench based on 5 different seeds. As shown in Fig. 9, the text-only model degrades sharply: –3.5% at 30 steps and –17% at 15 steps. The upsampled variant follows a similar trend but appears even more unstable (–10.2% and –22.0%), largely because it starts from a higher baseline at 50 steps. In contrast, the factorized model remains effectively unchanged, with only +1.2% and +0.3% variation. Although reducing diffusion from 50 to 15 steps would theoretically yield a 3.3× speedup, the full end-to-end pipeline—including anchor-image generation remains 2.1× faster in wall-clock time (1’39 vs. 3’30). Despite this substantial reduction in compute, the factorized model preserves the same accuracy, indicating that stronger conditioning not only improves quality but also stabilizes the diffusion trajectory, enabling far more efficient sampling.
50503030151530-3025-2520-2015-1510-105-5055Baseline (50 steps)Sampling stepsPercentage Change (%)$T2V$ (upsampled)Factorized T2V (Ours)
Figure 4: Percentage change in WAN 2.2 (5B) performance relative to the 50-step baseline. Both T2V variants degrade as steps decrease, while the factorized model remains stable even at 15 steps.
@@ 5.4 Impact on Diversity
We evaluate generation diversity under several experimental settings, with results reported in Tab. 4. For each variation we generate 25 videos per prompt and calculate the diversity metric introduced by Vbench2.0.
The standard diversity evaluation involves fixing a text prompt and varying the random seed. For a standard T2V model, this generates multiple videos from the same text. However, in our factorized pipeline, a single seed change affects both the anchor image generation (T2I stage) and the subsequent video diffusion model. To disentangle these two sources of variance, we define two settings. First, we vary the seed for the entire factorized pipeline. This captures the total diversity (variation in both scene and motion). This is reported as “Seeds” in Tab. 4. Second, we generate a single anchor image, fix it, and then vary the seed only for the video diffusin model. This isolates motion-level diversity. This is reported as “Single Image”. We observe that the full FVG pipeline yields diversity scores comparable to the baseline T2V model. Conversely, and as expected, fixing the anchor image and varying only the video model seed causes the diversity score to drop significantly. These experiments suggest that generation diversity could be usefully disentangled into scene-level diversity, which dominates the total variance, and motion-level diversity, which represents variations in animation for a static scene. Finally, we evaluate an additional diversity enhancement strategy by using an LLM to rephrase prompts. For each original prompt, we generate 25 semantic variations and provide them to the models. This “Prompt Rephrasing” method, as shown in Tab. 4, significantly boosts diversity scores across all tested models, confirming that semantic-level variation is a potent source of diverse outputs.

### v2

# Anchored Video Generation: Decoupling Scene Construction and Temporal Synthesis in Text-to-Video Diffusion Models

@@ Abstract
State-of-the-art Text-to-Video (T2V) diffusion models can generate visually impressive results, yet they still frequently fail to compose complex scenes or follow logical temporal instructions. In this paper, we argue that many errors, including apparent motion failures, originate from the model’s inability to construct a semantically correct or logically consistent initial frame. We introduce Anchored Video Generation (AVG), a modular pipeline that decouples these tasks by decomposing the Text-to-Video generation into three specialized stages: (1) Reasoning, where a Large Language Model (LLM) rewrites the video prompt to describe only the initial scene, resolving temporal ambiguities; (2) Composition, where a Text-to-Image (T2I) model synthesizes a high-quality, compositionally-correct anchor frame from this new prompt; and (3) Temporal Synthesis, where a video model, finetuned to understand this anchor, focuses its entire capacity on animating the scene and following the prompt. Our approach sets a new state-of-the-art on the T2V CompBench benchmark and significantly improves all tested models on VBench2. Furthermore, we show that visual anchoring allows us to cut the number of sampling steps by 70% without any loss in performance. AVG offers a simple yet practical path toward more efficient, robust, and controllable video synthesis. Project website: https://vita-epfl.github.io/AVG/
@@ 1 Introduction
Figure 1: Example failure modes in SoTA video generative models. We compare the first frame of videos from T2V Wan 2.2 [wan2025wan] and our Anchored Wan 2.2 under the AVG pipeline. T2V Wan 2.2 composes scenes incorrectly and exhibits logical temporal inconsistencies, struggling to establish coherent scene structure without explicit visual grounding.
@@
Our analysis across multiple state-of-the-art systems reveals that many errors originate in the model’s representation of the scene itself. In Figure˜1, we observe that existing models routinely (1) miscompose visual elements of the scene, and (2) generate initial states incompatible with the requested temporal evolution. Once the scene composition violates the prompt semantics, the temporal modeling fails. For example, if the model wrongly understands the context by assuming a “chicken" as a “roasted chicken" (see 1st sample), it becomes impossible to satisfy the prompt “a chicken tasting the sauce,” regardless of the strength of the motion model. This insight indicates a critical gap: Current T2V models may have over-entangled the two distinct tasks, static scene construction and temporal synthesis, yet are not structurally equipped to handle the first reliably.
@@
To delve into and understand this gap, we explore a diagnostic study by comparing Text-to-Video (T2V), Image-to-Video without text prompts (I2V), and Image-to-Video with text prompts (I2V+text) models in Table˜1, revealing two observations. T2V models frequently produce distributionally shifted frames yet still achieve FVD scores comparable to I2V models, indicating that their motion modeling remains reasonably natural even when scene fidelity is relatively poor. I2V models exhibit the complementary behavior, strong FID from accurate initial scenes and weaker temporal coherence, while I2V+text balances both aspects. This contrast suggests a structural mismatch in current T2V models: scene grounding and temporal synthesis are two distinct modeling challenges, yet existing architectures attempt to learn both simultaneously within a single model.
@@
These observations motivate us to revisit T2V. Rather than formulating it as a single task, we decompose it into two fundamentally sub-problems: specifying the initial scene and generating its temporal evolution. A natural way to operationalize this separation is through a visual anchor defining the intended starting state of the video. However, such an anchor cannot be generated by simply running a Text-to-Image (T2I) model on the original video prompt, which often contains multi-stage actions or temporal transformations. Obtaining an appropriate anchor requires first clarifying the prompt’s intended initial condition.
@@
Based on this insight, we propose Anchored Video Generation (AVG), a pipeline that explicitly decomposes T2V generation into specialized stages designed to address each subproblem independently: (1) Reasoning: A Large Language Model (LLM) rewrites the video prompt into an explicit description of the initial scene, removing temporal transitions and eliminating logical ambiguities. (2) Composition: A T2I model uses this refined prompt to synthesize a high-quality, semantically grounded anchor frame. (3) Temporal synthesis: A video diffusion model conditions on both the anchor and the original prompt, focusing exclusively on generating coherent temporal dynamics. We refer to this instantiated model variant as anchored T2V, as it augments a standard T2V backbone with an explicit intermediate visual anchor that establishes the intended scene composition. The contributions of this work are as follows:
@@
  • We introduce Anchored Video Generation, a simple and modular three-stage pipeline that separates reasoning, scene composition, and temporal synthesis, enabling improved scene grounding and temporal coherence.
@@
We will also release our models, code, benchmark prompts and image anchors generations to support reproducibility and facilitate further research.
@@ 2 Related Work
Spatial-Temporal Factorization. To address the complexity in video-centric learning, factorization [tran2018closer] has become a fundamental architectural strategy [liu2022video, bertasius2021space, lin2019tsm, ma2024latte], which decouples spatial and temporal components to enable efficient video representation. Some works [chen2023videocrafter1, videoworldsimulators2024] extend this paradigm via hierarchical factorization across multiple scales. At a higher level, Emu Video [girdhar2024factorizing] similarly explores factorizing text-to-video generation through full training with an explicit intermediate image. While Emu Video studies this idea through a dedicated factorized generator, our method is designed as a lightweight retrofit for existing pretrained T2V backbones, making the approach more modular and easier to integrate in practice. Existing architectural factorization approaches can also compromise spatial-temporal coupling under limited model capacity, leading to suboptimal results when data is scaled [ma2024latte]. Different from these lines of work, we lift factorization to a higher, task-oriented level, thereby preserving model capacity while achieving stronger text-following capabilities in video generation.
Figure 2: Overview of the finetuning and inference pipelines in AVG. Anchor-Grounding Finetuning: the Video Diffusion Model (VDM) is trained to follow a visual anchor by injecting a clean image latent at a randomly chosen frame position and setting its diffusion timestep to $t=0$. Lightweight LoRA finetuning teaches the model to treat this clean frame as a fixed scene constraint that guides the rest of the video. Anchored Generation Pipeline: an LLM rewrites the video prompt into a first-frame scene description to generate an anchor image. The anchor is then injected into the VDM with its timestep fixed to 0.
@@ 4 Methods
The goal of this work is to investigate how grounding text-to-video (T2V) diffusion models with at least one visual anchor affects their performance on complex compositional tasks. By introducing a visual anchor, we aim to decouple video generation into two subtasks, scene context construction and temporal motion modeling, and evaluate whether this separation improves the model’s understanding of scenes, objects, and their interactions. Our focus extends beyond visual quality to assess reasoning-oriented capabilities of video diffusion models. To this end, we compare standard T2V models with their text–image–to–video (I2V+text) counterparts across benchmarks designed to evaluate compositional video generation. Furthermore, to enable a fair comparison within the same architecture, we propose dedicated finetuning and inference pipelines that integrate a visual anchor into a pure T2V model. This allows the same backbone to operate either in a standard text-only mode or in an anchor-grounded mode. The following section details these pipelines which are summarized in Fig.˜2.
@@ 4.1 Anchor-Grounding Finetuning
A core capability required by the our approach is the ability to inject a visual anchor into a pretrained T2V diffusion model. We enable this through anchor-grounding finetuning, which teaches the model to treat the provided image as the initial frame of the generated video.
@@ 4.1.2 Training Procedure
The training procedure for adapting a T2V model to a visually anchored model is detailed in Algorithm 1. The core idea is to create a hybrid input $(z_{t},t_{vec})$ that pairs T1T-1 noisy frames at timestep $t$ with one clean anchor frame at $t=0$. The model $\epsilon_{\theta}$ is then finetuned via LoRA to denoise this out-of-distribution input, learning to treat the $t=0$ frame as a ground-truth anchor. The model remains text-conditioned through the original prompt $y$, while the visual anchor is introduced directly through the latent sequence.
@@
4.2 Anchored Generation Pipeline
At inference time, we first construct a visual anchor from the input text prompt before video sampling. Given a video text prompt $y$, we use a large language model (LLM) to generate a plausible detailed description of the scene’s initial state $y_{0}$, which emphasizes spatial layout, appearance, and object configuration. This reformulated prompt $y_{0}$ is then passed to a pretrained image generation model $\mathcal{G}_{\text{img}}$ to synthesize the anchor image $I_{a}=\mathcal{G}_{\text{img}}(y_{0})$.
@@ 5.2 Performance on Complex Tasks
Results on T2V-CompBench: As shown in Tab.˜2 across all models, adding an anchor image consistently improves compositional performances. All smaller anchored models (CogVideo 5B, Wan 5B and Wan 1B) outperform the larger Wan 14B T2V model. Our anchored Wan 5B also outperforms commercial PixVerse-V3 baseline which is the best reported model on the benchmark. This demonstrates that visual grounding substantially enhances scene and action understanding even in smaller-capacity models. Within each model family, the anchored version outperforms the original model. Notably, our lightweight anchor-grounded LoRA on WAN 14B reaches performance comparable its pretrained I2V 14B variant (0.661 vs. 0.666), despite requiring no full retraining.
@@
Anchored I2V + Text 0.927 0.217 0.680 0.320 0.850 0.886 0.781 0.666 (+21.09%)
Anchored T2V (Ours) 0.929 0.183 0.696 0.328 0.839 0.891 0.757 0.661 (+20.18%)
@@
Anchored T2V (Ours) 0.920 0.384 0.656 0.262 0.790 0.844 0.402 0.608 (+32.75%)
Results on VBench2.0: We summarize in Tab.˜3 the results on VBench 2.0, reporting the five aggregated metrics. The complete set of 18 underlying metrics is provided in the Appendix. We report only the composition score instead of the original creativity metric (the latter averages composition and diversity) and study the diversity specifically in Sec.˜5.4. Across all architectures, anchored T2V consistently improves every VBench metric except human fidelity, which shows a small decline even when using prompt upsampling. Similar to the observations on T2V-CompBench, our anchored Wan 5B outperforms the larger Wan 14B model, demonstrating that visual grounding is sometimes more beneficial than scaling. While performance gains are consistent, improvements on VBench are generally smaller than on T2V-CompBench. This is expected since VBench applies a stricter evaluation protocol as most metrics are binary (0 or 1 per video) with no partial credit. Nonetheless, the trend confirms that visual anchoring systematically enhances compositional reasoning and controllability even under stricter assessment criteria.
@@
Anchored T2V (Ours) 0.649 (+39.4%) 0.636 (+25.6%) 0.304 (+69.3%) 0.652 (–14.0%) 0.562 (+25.8%) 0.561 (+19.0%)
Qualitative Results. Figure˜5 shows representative examples comparing text-only generation with our anchored approach. Anchored videos consistently exhibit more accurate scene composition, stronger object–attribute binding, and clearer temporal progression.
Figure 3: Qualitative results showing our anchored method leads to better performance. Additional results are provided in the appendix.
@@ 5.3 Robustness to Reduced Sampling Steps
We assess robustness to reduced diffusion steps by decreasing sampling from 50 to 30 and 15 for Wan5B and calculate the results on T2V Compbench based on 5 different seeds. As shown in Fig.˜9, the text-only model degrades sharply: –3.5% at 30 steps and –17% at 15 steps. The upsampled variant follows a similar trend but appears even more unstable (–10.2% and –22.0%), largely because it starts from a higher baseline at 50 steps. In contrast, the anchored model remains effectively unchanged, with only +1.2% and +0.3% variation. Although reducing diffusion from 50 to 15 steps would theoretically yield a 3.3× speedup, the full end-to-end pipeline—including anchor-image generation remains 2.1× faster in wall-clock time (1’39 vs. 3’30). Despite this substantial reduction in compute, the anchored model preserves the same accuracy, indicating that stronger conditioning not only improves quality but also stabilizes the diffusion trajectory, enabling far more efficient sampling.
50503030151530-3025-2520-2015-1510-105-5055Baseline (50 steps)Sampling stepsPercentage Change (%)$T2V$ (upsampled)Anchored T2V (Ours)
Figure 4: Percentage change in WAN 2.2(5B) performance relative to 50-step baseline. Both T2V variants degrade as steps decrease, while ours remains stable at 15 steps.
@@ 5.4 Impact on Diversity
We evaluate generation diversity under several experimental settings, with results reported in Tab.˜4. For each variation we generate 25 videos per prompt and calculate the diversity metric introduced by Vbench2.0. The standard diversity evaluation involves fixing a text prompt and varying the seed. For a standard T2V model, this generates multiple videos from the same text. However, in our pipeline, a single seed change affects both the anchor image generation (T2I stage) and the subsequent video diffusion model. To disentangle these two sources of variance, we define two settings. First, we vary the seed for the entire pipeline. This captures the total diversity (variation in both scene and motion). This is reported as “Seeds" in Tab.˜4. Second, we generate a single anchor image, fix it, and then vary the seed only for the video diffusion model. This isolates motion-level diversity. This is reported as “Single Image". We observe that our full pipeline yields diversity scores comparable to the baseline T2V model. Conversely, and as expected, fixing the anchor image and varying only the video model seed causes the diversity score to drop significantly. These experiments suggest that generation diversity could be usefully disentangled into scene-level diversity, which dominates the total variance, and motion-level diversity, which represents variations in animation for a static scene. Finally, we evaluate an additional diversity enhancement strategy by using an LLM to rephrase prompts. For each original prompt, we generate 25 semantic variations and provide them to the models. This “Prompt Rephrasing" method, as shown in Tab.˜4, significantly boosts diversity scores across all tested models, confirming that semantic-level variation is a potent source of diverse outputs.